OpenStudy (anonymous):

What are the excluded values for x+4 10 ----- = ----- x−3 z+5 A. -3, 5 B. 3, -5 C. -4, 3, -5 D. There are no excluded values

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

What makes a value "excluded" @LaBlueEyes2013 ?

OpenStudy (skullpatrol):

$$\huge \frac{x+4}{x-3} =\frac{10}{z+5}$$